# GoalNet: Goal Areas Oriented Pedestrian Trajectory Prediction

### v1

@@ Abstract
Predicting the future trajectories of pedestrians on the road is an important task for autonomous driving. The pedestrian trajectory prediction is affected by scene paths, pedestrian’s intentions and decision-making, which is a multi-modal problem. Most recent studies use past trajectories to predict a variety of potential future trajectory distributions, which do not account for the scene context and pedestrian targets. Instead of predicting the future trajectory directly, we propose to use scene context and observed trajectory to predict the goal points first, and then reuse the goal points to predict the future trajectories. By leveraging the information from scene context and observed trajectory, the uncertainty can be limited to a few target areas, which represent the ”goals” of the pedestrians. In this paper, we propose GoalNet, a new trajectory prediction neural network based on the goal areas of a pedestrian. Our network can predict both pedestrian’s trajectories and bounding boxes. The overall model is efficient and modular, and its outputs can be changed according to the usage scenario. Experimental results show that GoalNet significantly improves the previous state-of-the-art performance by 48.7% on the JAAD and 40.8% on the PIE dataset.
1 Introduction
Autonomous driving is under rapid technological development and changes in recent years, and the first priority must be safety. In order to achieve this goal, one of the fundamental tasks is to predict the trajectories of pedestrians, which can help autonomous vehicles to avoid collisions in advance. For example, if we cannot predict that a pedestrian is about to cross the road, the vehicle can only start braking when the pedestrian appears in front of it. The short response time may cause failure to stop in time, resulting in casualties. Therefore, estimating pedestrian trajectories quickly and accurately is a crucial task. But trajectory prediction is not an easy task, after all, people’s actions and thoughts are constantly changing. There are two main factors that affect the choice of trajectory path. The first factor is the pedestrian’s past trajectory, which contains the implicit intention. The second factor is the surrounding environment, including passable areas and inaccessible areas. A good pedestrian trajectory prediction model needs to consider these two factors.
@@ 1 Introduction
The trajectory prediction part is usually regarded as a problem of time series analysis, so in the early processing of this problem, only the historical trajectory is used as input. However, if it is necessary to further consider the input of scene context at the same time, the two inputs of trajectory and environment need to be converted into the same dimension in the model, and the input of environment is a two-dimensional picture. Since the scene image is 2-dimensional, and the trajectory is 1-dimensional in the original data if the scene image is converted into the 1-dimensional alignment with the trajectory dimension, the spatial information will be lost, but if the trajectory is converted into the 2-dimensional heat map to align with the scene image, this problem can be solved.
In terms of trajectory prediction output, deterministic trajectory prediction was mostly carried out in the early days, while multi-modal trajectory prediction is the mainstream of current trajectory prediction research, that is, multi-modal prediction. Because the pedestrian will change the route of the next action due to different environmental conditions or his ideas at each time step, predicting multiple uncertain trajectories can better represent pedestrian’s future trajectory. Studies in [2, 5, 18, 32] also show that multi-modal trajectory prediction is more accurate than deterministic trajectory prediction, and SGNet [34] has also achieved remarkable achievements in uncertainty prediction.
Unlike the regression problem, which treats trajectories directly as sequential patterns, some previous studies, such as Y-Net [25], proposed to address the problem in a novel way. The authors believe that pedestrian trajectory prediction is different from normal linear trajectory because the decision-making subject of trajectory is human, and it is obviously different from general time series forecasting because there are more potential variables of uncertainty in decision-making. The authors of Y-Net mentioned that such uncertain decision factors would have a greater impact on long-term trajectory prediction, and propose a long-term trajectory prediction network with good performance. Our work is inspired by Y-Net, but different from Y-Net’s fixed BEV perspective. In our work, we make trajectory prediction based on vehicle FPV perspective, which indicates that the decision variables of the trajectory besides the pedestrians, the driver’s decision will have a similar impact on the future movement of the trajectory. And the speed of the vehicle will make these potential decision factors further amplified. Therefore, due to the more complex environmental conditions and the influence of more uncertain factors, the uncertainty trajectory prediction has become more important in our work.
@@
2 Related Work
Nowadays, in the field of autonomous driving, the datasets used can mainly divided into two types. One is the datasets containing LiDAR data, such as nuScenes [4] or Argoverse[6]. After using LiDAR data, the dataset can be divided into two types. The ability to 3D detect and track the detected object and to measure its velocity and acceleration makes it possible to increase the accuracy of training autonomous driving-related neural networks by adding effective auxiliary information. In this dataset, some recent studies such as ViP3D [11] have achieved remarkable results. The other is a dataset with only camera images, such as JAAD and PIE, which are cheaper, smaller, and easier to install on vehicles than LiDAR. Especially in recent years, most vehicles have installed dashboard cameras. This means that the trajectory prediction model designed based on the scenarios of such datasets has a broader and universal application prospect. Moreover, the neural network does not need to process the LiDAR data, the memory resources used are less and the execution speed is more advantageous. For most vehicle safety assistance systems with limited hardware resources, these are the characteristics of such embedded systems with extremely limited computing resources.
@@ 2 Related Work
Bird’s Eye View (BEV).
Pedestrian trajectory prediction on BEV mainly simplifies the factor of self-motion, and does the trajectory prediction under the situation of fixed viewing angle and position, and the common datasets are ETH/UCY and SDD. In ETH/UCY, the Y-Net [25] proposed by Mangalam et al. contains RGB scenes and historical trajectories as inputs, semantic segmentation of the scenes and estimation of trajectory endpoint locations and the probability distribution of the path points in them are used for trajectory prediction;The $V^{2}$-Net [35] proposed by Wong et al. adds Fourier transformations to model and predict the trajectories in the spectral trajectory view, a kind of Transformer-based network containing coarse-level keypoints estimation and fine-level spectrum interpolation, this has led to an emphasis on pedestrian socialization and interaction with the scene; The NSP-SFM [39] proposed by Yue et al. uses a deterministic model based on microscopic equations, and the key parameters can be learned from the data and integrated into the NSP model using an autoencoder that captures the kinematic dynamics and uncertainty. Trajectory prediction in a top-down view such as the Standford Drone Dataset has also yielded better results in the $V^{2}$-Net model and NSP-SFM mentioned above.In addition, the TDOR [12] proposed by Guo et al. establishes the CVAE by taking the future trajectories in the neighboring grids as interpretable potential variables. They are iterated in maximum-entropy inverse reinforcement learning through differentiable value iteration networks and to maximize the data possibilities for learning path planning and trajectory generation.
@@
First Person View (FPV).
Pedestrian trajectory prediction with a first person perspective is more suitable for training automated driving systems, because it includes the vehicle’s self-motion, in which JAAD and PIE datasets are more widely used, and we also used these two datasets for the validation. In 2019 the PIE dataset was released by Rasouli et al. They also proposed an intention estimation and trajectory prediction model with a large number of LSTM modules [28], which achieved good results not only on its own dataset but also on JAAD. Yao et al. proposed the Bitrap model [38], this trajectory prediction model consists of four parts: the condition prior network, the identification network, the target generation network and the trajectory generation network, and uses a large number of GRU modules in the model, It is a multi-modal trajectory prediction model that uses bounding boxes or trajectories as inputs and outputs multiple predicted trajectories. Recently Wang et al. indicated that human intention is a manifestation of planning, and intention at previous time points also changes their view in the present, and proposed SGNet [34], which combines past goals into an encoder through average pooling, and uses an attention mechanism to learn the importance of each goal to the prediction. And both Bitrap and SGNet perform well on JAAD and PIE dataset.
@@
3 Proposed Method
The problem of trajectory prediction can be defined as follows. Suppose that the current time is t, and P represents the bounding box coordinates of the object with length and width (in pixels), then Pt represents the position of the object at time t. The given observation time is 15 frames, X = {Pt-14, Pt-13, …, Pt}, our goal is to predict the position of X in the next 45 frames, Y = {Pt+1, Pt+2, …, Pt+45}. For the prediction of stochastic trajectories, given an observation trajectory X, there may be multiple results, so various trajectories corresponding to these uncertain futures need to be generated to cover this uncertainty, the uncertain trajectories are expressed as S, and the corresponding n trajectories are generated. S = {S1, S2, …, Sn}.
GoalNet consists of three modules, which are feature extraction, trajectory prediction, and bounding box prediction module, and we estimate the number of parameters and FLOPs of our model as shown in Table 1.
@@ 3 Proposed Method
Table 1: GoalNet #param. & FLOPs.
3.1 Feature Extract Module
Scene context is very important for trajectory prediction, but many previous studies[1, 38, 34, 13] only deal with trajectory without considering environmental images. On specific trajectory datasets (e.g., Standford Drone Dataset (SDD) [30]), some studies [22, 21] begin to use segmentation as an environmental feature extraction method. However, for most trajectory prediction datasets, In particular, it is not an easy task to construct segmentation annotation for first person perspective (FPV) datasets, and it is difficult to define appropriate classes, and the output channels are not very flexible, because they depend on the number of segmentation classes. If RGB scene images are directly input into the trajectory prediction model, the trajectory prediction model needs to adapt to the processing of environmental images and trajectory information at the same time. In order to make the trajectory prediction model consider both scene context and trajectory information, the trajectory prediction module focuses on trajectory coding. Rather than directly input RGB images into the trajectory prediction model, it is necessary to establish the image and maintain the hidden state of complete spatial information through feature extraction. Indirectly inputting the scene image into the trajectory prediction module is a better method.
@@ 3.2 Trajectory Module
This module encodes the observed trajectory and the scene image, and the decoder predicts the goal areas represented by a probability distribution, and calculates the main goal through spatial soft arg-max [9]. At the same time, K-means clustering algorithm is used to calculate the secondary goals in the goal areas, and these goals are input back to the decoder in sequence again to generate uncertainty trajectories with these goals as future assumptions.
@@
Since trajectory prediction does not require a special ability to perceive the edge of an object like segmentation, stride 2 convolution, like modern convolutional networks such as ResNet [15] and ConvNeXt [23] is probably more advantageous in this case. Therefore, we replace the U-Net downsample method with stride 2 convolution. At the same time, by replacing the convolution layer with ASPP at the deepest (Center layer) of the trajectory prediction module, the perception field of view of the model is greatly improved, and the spatial information of the trajectory is more fully utilized.
@@
3.3 Bounding Box Module
In Figure 4, in order to predict the bounding box corresponding to the trajectory, we convert the center point trajectory predicted by the trajectory prediction module to the trajectory in the form of bounding box by predicting the width and height of the bounding box. Different from previous studies [38, 34], previous studies directly predicted bounding box location, coordinates sequences representing the upper left and lower right points of bounding box, and then calculated the trajectory of its center point. In our work, it is not easy to generate two points at a time in a single time step, so we adopt the opposite method, that is, first predict the location, coordinates of the center point, and then predict the trajectory in the form of bounding box. As shown in the table 1, the computational consumption of this method can be almost ignored.
@@ 3.3 Bounding Box Module
3.4 Loss Functions
Since the predicted trajectory is the probability distribution of each timestep, and the feature extraction module does not make an actual prediction, but only generates environmental features for the trajectory prediction module. Therefore, we use binary cross-entropy loss to supervise the prediction of trajectory feature extraction and prediction module. For the bounding box prediction module, since the prediction results are absolute values of width and height, we use SmoothL1Loss [10] to monitor. Thus, for each training sample, our final loss is summarized as follows,
@@ 4 Experiments
Table 2: Stochastic results (K=20) on JAAD and PIE in terms of MSE/CMSE/CFMSE. ↓ denotes lower is better.
@@
Table 3: Exploration study of our model on JAAD. ↓ denotes lower is better. Our final proposal is marked in gray.
@@
We use ConvNeXt-Tiny [23] as the backbone of the feature extraction module, convert the two-dimensional input of RGB 3 channels into two-dimensional output of 6 channels, and implement the encoder-decoder of the trajectory prediction module using a U-Net-like structure. The input of the trajectory prediction module is taken by the observed 15-time step trajectory heatmap as 15 channels and concatenate with 6 channels scene feature map. The channel size of the encoder is 32-32-64-64-64-128, while the decoder is the opposite, and the trajectory prediction output of 45 time steps is generated. The last point of the first output trajectory is feedback to the decoder as a goal to generate the final future trajectories. All models were trained with a batch size of 8, the Adam [19] optimizer with an initial learning rate of 1e-4, and the ReduceLROnPlateau scheduler on a single RTX2080Ti GPU.
@@
4.1 Experiments on JAAD and PIE datasets
We compared the results with the following baseline models: 1) BiTraP-GMM 2) BiTraP-NP 3) SGNet-ED. To make a fair comparison of stochastic results with [38, 34] in JAAD and PIE Benchmarks, we generate K = 20 proposals and report the best sample. Table 2 shows the trajectory prediction results of JAAD and PIE datasets. The results show that our method significantly outperforms previous state-of-the-art methods by an average of 48.7% on JAAD and 40.8% on PIE.
4.2 Exploration Study
@@
To design our GoalNet model and make predictions on the JAAD and PIE datasets, we use the U-Net [31] architecture as a base, and reference modern convolutional neural network designs such as ConvNeXt [23]. Evaluate the submodules and micro-designs of the overall GoalNet design, the result is shown in Table 3.
Table 2(a) analyzes the impact of removing different modules on GoalNet. The results show that the feature extraction module is necessary. If the environment RGB image is directly input into the trajectory prediction module, the final accuracy will be greatly reduced. And removing either the Attention Gate (AG) [26] or Atrous Spatial Pyramid Pooling (ASPP) [7] modules would result in lower final accuracy, so adding AG and ASPP does have a positive effect on the model.
Table 2(b) analyzes the performance using different normalization layers, and while Batch Normalization (BN) [17] is an important and common component in convolutional neural networks and is preferred for many visual tasks, because it improves the convergence of the model and reduces overfitting, but experiments show that BN is very unsuitable for use in our structure, and the experimental data show that Group Normalization (GN) [36] is better than BN. However, the implementation of layer normalization of GN is very similar to BN, the difference is that GN avoids some of the batch restrictions of BN by group normalization of channels within a single training sample, and we speculate that normalization on batches will destroy some important information in the training data. In recent years, Layer Normalization (LN) [3]has gradually increased in importance and replaced BN gradually, including replacing BN with LN in transformer and ConvNeXt. ConvNeXt shows that it is not difficult to replace BN with LN in convolutional neural networks, and it is possible to achieve better results. Our experiments also show that LN does perform best, so LN is used as the layer normalization choice in our model.
The results of different activation functions are discussed in Table 2(c). Rectified Linear Unit (ReLU) is widely used in convolutional neural networks, and due to the dying ReLU problem, ReLU has therefore derived many variants (e.g., LeaklyReLU, PReLU, GELU) to solve this problem. Experiments have shown that ReLU performs best in our model, so we have chosen to continue using ReLU in our model.
The influence of different downsampling layers is studied in Table 2(d). Max Pooling is often used in previous networks, it retains key features by taking the maximum value during downsampling. In addition, Max Pooling’s gradient flow propagation is easier than stride 2 convolution, which makes it easier for neural networks to learn. However, in modern convolutional neural networks, stride 2 convolution is often used to replace Pooling, because stride 2 convolution has learnable parameters. With appropriate design, some accuracy advantages may be achieved.
@@
4.3 Qualitative Results
Figure 6 and Figure 7 show GoalNet’s prediction results on the test split of the JAAD and PIE datasets, respectively. (a) Show the prediction results for all 20 uncertain trajectories. Most of the multi-modal prediction results are close to the ground truth trajectory in the future, and the high probability distribution near the ground truth trajectory is successfully predicted, which also indicates the stability of the model and converges the uncertainty to a certain range. (b) The best bounding box trajectory is shown. The results show that our model can correctly predict the scaling of the bounding box as the pedestrian approaches or moves away from the vehicle.

### v2

@@ Abstract
Predicting the future trajectories of pedestrians on the road is an important task for autonomous driving. The pedestrian trajectory prediction is affected by scene paths, pedestrian’s intentions and decision-making, which is a multi-modal problem. Relying solely on historical coordinates pedestrian data represents the most straightforward method for pedestrian trajectory prediction. Nevertheless, the accuracy achieved by this method is limited, primarily because it fails to account for the crucial scene paths impacting the pedestrian. Instead of predicting the future trajectory directly, we propose to use scene context and observed trajectory to predict the goal points first, and then reuse the goal points to predict the future trajectories. By leveraging the information from scene context and observed trajectory, the uncertainty can be limited to a few target areas, which represent the ”goals” of the pedestrians. In this paper, we propose GoalNet, a new trajectory prediction neural network based on the goal areas of a pedestrian. Our network can predict both pedestrian’s trajectories and bounding boxes. The overall model is efficient and modular, and its outputs can be changed according to the usage scenario. Experimental results show that GoalNet significantly improves the previous state-of-the-art performance by 48.7% on the JAAD and 40.8% on the PIE dataset.
Index Terms:
pedestrians, trajectory prediction, future trajectories.
[…]
=-21pt
I Introduction
Trajectory prediction offers significant advantages across various domains by enabling proactive decision-making and enhancing safety. For example, in autonomous driving navigation, it is crucial for anticipating the movements of pedestrians and other vehicles, thereby facilitating collision avoidance and optimizing path planning. similarly, in vessel dynamic positioning [1], accurate trajectory forecasting supports precise station-keeping and safe maneuvering within complex maritime environments. Autonomous driving is under rapid technological development and changes in recent years, and the first priority must be safety. In order to achieve this goal, one of the fundamental tasks is to predict the trajectories of pedestrians, which can help autonomous vehicles to avoid collisions in advance. For example, if we cannot predict that a pedestrian is about to cross the road, the vehicle can only start braking when the pedestrian appears in front of it. The short response time may cause failure to stop in time, resulting in casualties. Therefore, estimating pedestrian trajectories quickly and accurately is a crucial task. But trajectory prediction is not an easy task, after all, people’s actions and thoughts are constantly changing. There are two main factors that affect the choice of trajectory path. The first factor is the pedestrian’s past trajectory, which contains the implicit intention. The second factor is the surrounding environment, including passable areas and inaccessible areas. A good pedestrian trajectory prediction model needs to consider these two factors.
@@ I Introduction
Trajectory prediction is frequently conceptualized as a time series analysis problem, which led early approaches to primarily rely on historical trajectories for input. However, for the purpose of trajectory prediction utilizing a full CNN architecture, it is essential to additionally consider the incorporation of scene context. This requires that the trajectory input and the two-dimensional environmental image be transformed to achieve a consistent dimensionality within the model’s framework. Since the scene image is 2-dimensional, and the trajectory is 1-dimensional in the original data if the scene image is converted into the 1-dimensional alignment with the trajectory dimension, the spatial information will be lost, but if the trajectory is converted into the 2-dimensional heat map to align with the scene image, this problem can be solved.
In terms of trajectory prediction output, deterministic trajectory prediction was mostly carried out in the early days, while multi-modal trajectory prediction is the mainstream of current trajectory prediction research. Because the pedestrian will change the route of the next action due to different environmental conditions or his ideas at each time step, predicting multiple uncertain trajectories can better represent pedestrian’s future trajectory. Studies in [11, 12, 13, 14] also show that multi-modal trajectory prediction is more accurate than deterministic trajectory prediction, and SGNet [7] has also achieved remarkable achievements in uncertainty prediction.
Unlike the regression problem, which treats trajectories directly as sequential patterns, some previous studies, such as Y-Net [5], proposed to address the problem in a novel way. The authors argue that pedestrian trajectory prediction is different from normal linear trajectory because the decision-making subject of trajectory is human, and it is obviously different from general time series forecasting because there are more potential variables of uncertainty in decision-making. The authors of Y-Net mentioned that such uncertain decision factors would have a greater impact on long-term trajectory prediction, and propose a long-term trajectory prediction network with good performance. Our work is inspired by Y-Net, but different from Y-Net’s fixed BEV perspective. In our work, we make trajectory prediction based on vehicle first person view (FPV) perspective, which indicates that the decision variables of the trajectory besides the pedestrians, the driver’s decision will have a similar impact on the future movement of the trajectory. And the speed of the vehicle will make these potential decision factors further amplified. Therefore, due to the more complex environmental conditions and the influence of more uncertain factors, the uncertainty trajectory prediction has become more important in our work.
@@
II Related Work
Nowadays, in the field of autonomous driving, the datasets used can mainly divided into two types. One is the datasets containing LiDAR data, such as nuScenes [34] or Argoverse[35]. After using LiDAR data, the dataset can be divided into two types. The ability to 3D detect and track the detected object and to measure its velocity and acceleration makes it possible to increase the accuracy of training autonomous driving-related neural networks by adding effective auxiliary information. Some recent studies such as ViP3D [36] have achieved remarkable results on nuScenes dataset. The other is a dataset with only camera images, such as JAAD and PIE, which are cheaper, smaller, and easier to install on vehicles than LiDAR. Especially in recent years, most vehicles have installed dashboard cameras. This means that the trajectory prediction model designed based on the scenarios of such datasets has a broader and universal application prospect. When the neural network is designed to forego the processing of LiDAR data, it consequently reduces memory resource consumption and offers improved computational efficiency during execution. For most vehicle safety assistance systems with limited hardware resources, these are the characteristics of such embedded systems with extremely limited computing resources.
@@ II Related Work
II-A Bird’s Eye View (BEV).
Predicting pedestrian trajectories using a static single camera’s Bird’s-Eye View (BEV) primarily simplifies the challenge of accounting for self-motion. In this setup, trajectory prediction happens from a fixed viewpoint and position. Common datasets for this approach include ETH/UCY and SDD. In ETH/UCY, the Y-Net [5] proposed by Mangalam et al. contains RGB scenes and historical trajectories as inputs, semantic segmentation of the scenes and estimation of trajectory endpoint locations and the probability distribution of the path points in them are used for trajectory prediction;The $V^{2}$-Net [16] proposed by Wong et al. adds Fourier transformations to model and predict the trajectories in the spectral trajectory view, a kind of Transformer-based network containing coarse-level keypoints estimation and fine-level spectrum interpolation, this has led to an emphasis on pedestrian socialization and interaction with the scene; Goal-driven Long-Term Trajectory Prediction [2] proposed by Chen et al. introduces a method that enhances long-term forecasting by explicitly predicting a set of likely future goals for agents, subsequently generating diverse trajectories conditioned on these predicted goals. Its specialty lies in a dual-channel neural network architecture—comprising a Goal channel and a Trajectory channel—which employs Gated Recurrent Units (GRUs) to model temporal dependencies, automatically identifies and ranks destinations within the scene, and allows for generalization and transfer learning to unseen environments; The NSP-SFM [17] proposed by Yue et al. uses a deterministic model based on microscopic equations, and the key parameters can be learned from the data and integrated into the NSP model using an autoencoder that captures the kinematic dynamics and uncertainty. Trajectory prediction in a top-down view such as the Standford Drone Dataset has also yielded better results in the $V^{2}$-Net model and NSP-SFM mentioned above. In addition, the TDOR [18] proposed by Guo et al. establishes the CVAE by taking the future trajectories in the neighboring grids as interpretable potential variables. They are iterated in maximum-entropy inverse reinforcement learning through differentiable value iteration networks and to maximize the data possibilities for learning path planning and trajectory generation. The PPNet introduce a modular framework named PPNet for predicting pedestrian trajectories using multiple modalities. To improve prediction accuracy, PPNet incorporates a goal generator that models various distributions of potential goals, allowing it to identify and categorize diverse latent intentions. With the guidance of these goals, the anchor generator utilizes Transformer encoder-decoder networks to create anchors. Ultimately, the proposal generator combines the current position, anchors, and goals to produce a set of smooth, efficient, and probabilistic proposals. [44] The combination of YOLO, DeepSORT, and LSTM effectively addresses the issues of occlusion and interactions between pedestrians, which are primary causes of errors in trajectory prediction. YOLO is used for accurate pedestrian detection, while DeepSORT helps with robust tracking by maintaining consistent identities even when pedestrians overlap or move in close proximity. Meanwhile, LSTM models integrated dot product attention mechanism with with the temporal dependencies in the trajectories, allowing the system to predict future positions more accurately despite these challenges. This integrated approach helps minimize errors caused by occlusions and interactions, improving overall trajectory prediction performance. [47]
@@
II-B First Person View (FPV)
Pedestrian trajectory prediction with a first person perspective is more suitable for training automated driving systems, because it includes the vehicle’s self-motion, in which JAAD and PIE datasets are more widely used, and we also used these two datasets for the validation. In 2019 the PIE dataset was released by Rasouli et al. They also proposed an intention estimation and trajectory prediction model with a large number of LSTM modules [4], which achieved good results not only on its own dataset but also on JAAD. Yao et al. proposed the BiTraP model [6], this trajectory prediction model consists of four parts: the condition prior network, the identification network, the target generation network and the trajectory generation network, and uses a large number of GRU modules in the model, It is a multi-modal trajectory prediction model that uses bounding boxes or trajectories as inputs and outputs multiple predicted trajectories. Recently Wang et al. indicated that human intention is a manifestation of planning, and intention at previous time points also changes their view in the present, and proposed SGNet [7], which combines past goals into an encoder through average pooling, and uses an attention mechanism to learn the importance of each goal to the prediction. And both Bitrap and SGNet perform well on JAAD and PIE dataset. Several methods emphasize the pose representation of pedestrians. Zhou et al. integrated Bayesian Networks (BN) and SWI-PROLOG with the knowledge graph technique to infer three types of intentions that influence pedestrian crossing trajectory prediction. Subsequently, pedestrian intentions, historical trajectories, and body posture angle features were combined for trajectory prediction using a Bi-LSTM+Attention network. The enhanced Casrel model, along with pedestrian crossing intention estimation and trajectory prediction, was validated through experimental evaluation.The experiments demonstrated that the Casrel model successfully extracted triplet information, while the BN intention inference model effectively predicted pedestrian crossing intentions. [48]
@@ II-B First Person View (FPV)
III Proposed Methods
The problem of trajectory prediction can be defined as follows. Suppose that the current time is t, and P represents the bounding box coordinates of the object with length and width (in pixels), then Pt represents the position of the object at time t. The given observation time is 15 frames or 0.5 seconds in 30 frames per seconds (FPS) video, X = {Pt-14, Pt-13, …, Pt}, our goal is to predict the position of X in the next 45 frames or 1.5 seconds, Y = {Pt+1, Pt+2, …, Pt+45}. For the prediction of stochastic trajectories, given an observation trajectory X, there may be multiple results, so various trajectories corresponding to these uncertain futures need to be generated to cover this uncertainty, the multi-modal trajectories are expressed as S, and the corresponding n trajectories are generated. S = {S1, S2, …, Sn}.
GoalNet consists of three modules, which are feature extraction, trajectory prediction, and bounding box prediction module, and we estimate the number of parameters and FLOPs of our model as shown in Table I.
@@ III Proposed Methods
TABLE I: GoalNet #param. & FLOPs.
III-A Feature Extract Module
Before scene context was widely used as important data for trajectory prediction, some previous studies [19, 6, 7, 20] dealt with trajectories without incorporating environmental images. On specific trajectory datasets (e.g., Standford Drone Dataset (SDD) [10]), some studies [21, 22] begin to use segmentation as an environmental feature extraction method. However, for most trajectory prediction datasets, In particular, it is not an easy task to construct segmentation annotation for first person perspective (FPV) datasets, it is difficult to define appropriate classes, and the output channels are not very flexible, because they depend on the number of segmentation classes. If RGB scene images are directly input into the trajectory prediction model, the trajectory prediction model needs to adapt to the processing of environmental images and trajectory information at the same time. In order to make the trajectory prediction model consider both scene context and trajectory information, the trajectory prediction module focuses on trajectory coding. Rather than directly input RGB images into the trajectory prediction model, it is necessary to establish the image and maintain the hidden state of complete spatial information through feature extraction. Indirectly inputting the scene image into the trajectory prediction module is a better method.
@@ III-B Trajectory Module
This module encodes the observed trajectory and the scene image, and the decoder predicts the goal areas represented by a probability distribution, and calculates the main goal through spatial soft arg-max [23]. At the same time, K-means clustering algorithm is used to calculate the secondary goals in the goal areas, and these goals are input back to the decoder in sequence again to generate multi-modal trajectories with these goals as future assumptions.
@@
Since trajectory prediction does not require a special ability to perceive the edge of an object like segmentation, stride 2 convolution, like modern convolutional networks such as ResNet [29] and ConvNeXt [15] is probably more advantageous in this case. Therefore, we replace the U-Net downsample method with stride 2 convolution. At the same time, by replacing the convolution layer with ASPP at the deepest (Center layer) of the trajectory prediction module, receptive field of the model is greatly improved, and the spatial information of the trajectory is more fully utilized.
@@
III-C Bounding Box Module
In Figure 4, in order to predict the bounding box corresponding to the trajectory, we convert the center point trajectory predicted by the trajectory prediction module to the trajectory in the form of bounding box by predicting the width and height of the bounding box. Different from previous studies [6, 7], previous studies directly predicted bounding box location, coordinates sequences representing the upper left and lower right points of bounding box, and then calculated the trajectory of its center point. In our work, it is not easy to generate two points at a time in a single time step, so we adopt the opposite method, that is, first predict the location, coordinates of the center point, and then predict the trajectory in the form of bounding box. As shown in the table I, the computational consumption of this method can be almost ignored.
@@ III-C Bounding Box Module
III-D Loss Functions
The final layer of the trajectory module generates a heatmap of probability distribution, which is localized by spatial soft arg-max. Consequently, binary cross-entropy loss is applied to supervise the module. For the bounding box prediction module, since the prediction results are absolute values of width and height, we use SmoothL1 loss [33] to monitor. Thus, for each training sample, our final loss is summarized as follows,
@@ IV Experiments
TABLE II: Stochastic results (K=20) on JAAD and PIE in terms of MSE/CMSE/CFMSE. ↓ denotes lower is better.
@@
TABLE III: Exploration study of our model on JAAD. ↓ denotes lower is better. Our final proposal is marked in gray.
@@
We use ConvNeXt-Tiny [15] as the backbone of the feature extraction module, convert the two-dimensional input of RGB 3 channels into two-dimensional output of 6 channels, and implement the encoder-decoder of the trajectory prediction module using a U-Net-like structure. The input of the trajectory prediction module is taken by the observed 15-time step trajectory heatmap as 15 channels and concatenate with 6 channels scene feature map. The channel size of the encoder is 32-32-64-64-64-128, while the decoder is the opposite, and the trajectory prediction output of 45 time steps is generated. The last point of the first output trajectory is feedback to the decoder as a goal to generate the final future trajectories. All models were trained on a single RTX2080Ti GPU using a batch size of 8, the Adam [39] optimizer with an initial learning rate of 1e-4, and the ReduceLROnPlateau scheduler.
@@
IV-A Experiments on JAAD and PIE datasets
We compared the results with the following baseline models: 1) BiTraP-GMM 2) BiTraP-NP 3) SGNet-ED. To ensure a fair comparison with the stochastic results detailed in [6] and [7] for the JAAD and PIE Benchmarks, we select and present the most optimal trajectory from a collection of K=20 multi-modal generative trajectories. Table II shows the trajectory prediction results of JAAD and PIE datasets. The results show that our method significantly outperforms previous state-of-the-art methods by an average of 48.7% on JAAD and 40.8% on PIE.
IV-B Exploration Study
To design our GoalNet model and make predictions on the JAAD and PIE datasets, we use the U-Net [24] architecture as a base, and reference modern convolutional neural network designs such as ConvNeXt [15]. Evaluate the submodules and micro-designs of the overall GoalNet design, the result is shown in Table III.
Table III(a) analyzes the impact of removing different modules on GoalNet. The results show that the feature extraction module is necessary. If the environment RGB image is directly input into the trajectory prediction module, the final accuracy will be greatly reduced. And removing either the Attention Gate (AG) [30] or Atrous Spatial Pyramid Pooling (ASPP) [28] modules would result in lower final accuracy, so adding AG and ASPP does have a positive effect on the model.
Table III(b) analyzes the performance using different normalization layers, and while Batch Normalization (BN) [25] is an important and common component in convolutional neural networks and is preferred for many visual tasks, because it improves the convergence of the model and reduces overfitting, but experiments show that BN is very unsuitable for use in our structure, and the experimental data show that Group Normalization (GN) [26] is better than BN. However, the implementation of layer normalization of GN is very similar to BN, the difference is that GN avoids some of the batch restrictions of BN by group normalization of channels within a single training sample, and we speculate that normalization on batches will destroy some important information in the training data. In recent years, Layer Normalization (LN) [27] has gradually increased in importance and replaced BN gradually, including replacing BN with LN in transformer and ConvNeXt. ConvNeXt shows that it is not difficult to replace BN with LN in convolutional neural networks, and it is possible to achieve better results. Our experiments also show that LN does perform best, so LN is used as the layer normalization choice in our model.
The results of different activation functions are discussed in Table III(c). Rectified Linear Unit (ReLU) is widely used in convolutional neural networks, and due to the dying ReLU problem, ReLU has therefore derived many variants (e.g., LeakyReLU, PReLU, GELU) to solve this problem. Experiments have shown that ReLU performs best in our model, so we have chosen to continue using ReLU in our model.
The influence of different downsampling layers is studied in Table III(d). Max Pooling is often used in previous networks, it retains key features by taking the maximum value during downsampling. In addition, Max Pooling’s gradient flow propagation is easier than stride 2 convolution, which makes it easier for neural networks to learn. However, in modern convolutional neural networks, stride 2 convolution is often used to replace Pooling, because stride 2 convolution has learnable parameters. With appropriate design, some accuracy advantages may be achieved.
IV-C Qualitative Results
Figure 6 and Figure 7 show GoalNet’s prediction results on the test split of the JAAD and PIE datasets, respectively. (a) Show the prediction results for all 20 multi-modal trajectories. Most of the multi-modal prediction results are close to the ground truth trajectory in the future, and the high probability distribution near the ground truth trajectory is successfully predicted, which also indicates the stability of the model and converges the uncertainty to a certain range. (b) The best bounding box trajectory is shown. The results show that our model can correctly predict the scaling of the bounding box as the pedestrian approaches or moves away from the vehicle.